What is the reactance of a 5.2 microF capacitor at a frequency of 300Hz?
Answer #1

Capacitive reactance XC = 1/WC

Xc = 1/(2pifC)

Xc = 1/(2 *3.14 * 300 * 5.2 *10^-6)

Xc = 102.07 ohms